### v1

# SNE-RoadSegV2: Advancing Heterogeneous Feature Fusion and Fallibility Awareness for Freespace Detection

Abstract
Feature-fusion networks with duplex encoders have proven to be an effective technique to solve the freespace detection problem. However, despite the compelling results achieved by previous research efforts, the exploration of adequate and discriminative heterogeneous feature fusion, as well as the development of fallibility-aware loss functions remains relatively scarce. This paper makes several significant contributions to address these limitations: (1) It presents a novel heterogeneous feature fusion block, comprising a holistic attention module, a heterogeneous feature contrast descriptor, and an affinity-weighted feature recalibrator, enabling a more in-depth exploitation of the inherent characteristics of the extracted features, (2) it incorporates both inter-scale and intra-scale skip connections into the decoder architecture while eliminating redundant ones, leading to both improved accuracy and computational efficiency, and (3) it introduces two fallibility-aware loss functions that separately focus on semantic-transition and depth-inconsistent regions, collectively contributing to greater supervision during model training. Our proposed heterogeneous feature fusion network (SNE-RoadSegV2), which incorporates all these innovative components, demonstrates superior performance in comparison to all other freespace detection algorithms across multiple public datasets. Notably, it ranks the 1st on the official KITTI Road benchmark.
Index Terms:
Freespace detection, Feature fusion, semantic segmentation, deep learning.I Introduction
As a vital piece of the autonomous driving puzzle, reliable collision-free space (freespace for short) detection holds significant importance in autonomous driving systems, as it directly impacts a vehicle’s ability to make informed decisions and ensure dependable navigation [1]. In recent years, freespace detection has attracted considerable attention in research, with ongoing efforts aimed at addressing corner cases within complex and dynamic environments. Nevertheless, regardless of whether the approach is explicit programming-based or data-driven, the utilization of 3D information is growing in significance for freespace detection, primarily due to the valuable spatial geometry information it provides [2]. Feature-fusion networks with duplex encoders, designed to extract heterogeneous features from multiple data sources or modalities and fuse them to provide a more comprehensive understanding of the environment, have emerged as a viable solution to tackle this problem [3, 4].
The performance of a feature-fusion freespace detection network depends not only on the input data type but also on the manner in which these features are fused [5, 6]. A current bottleneck lies in the simplistic and indiscriminate fusion of heterogeneous features, often causing conflicting feature representations and erroneous detection results [2]. Taking the SNE-RoadSeg series [1, 7] as an example, their adopted feature fusion strategy essentially performs an element-wise summation between RGB and surface normal feature maps at each stage, neglecting the inherent differences in feature characteristics and their respective reliability [8]. Furthermore, as the network goes deeper, such an asymmetric feature fusion strategy tends to diminish the proportion of RGB features in the decoder’s input. This, in turn, leads to unsatisfactory performance, particularly in areas such as pavements or other lanes, where surface normals closely resemble those of freespace, demanding a greater reliance on color or textural information.
In the decoder aspect, we observe two phenomena: (1) inter-scale skip connections provide an advantage in achieving more comprehensive feature decoding, primarily due to their capability to capture both fine-grained and coarse-grained details [9], but unfortunately they are underutilized; (2) the currently adopted intra-scale skip connections appear to be excessively redundant for this task. Moreover, the utilization of pixel-wise binary cross-entropy (BCE) loss has been a common practice in freespace detection. It is, nevertheless, regrettable that no prior endeavors have been undertaken to delve into the fallible cases, notably the misclassifications occurring near semantic-transition and depth-inconsistent regions.
To address the aforementioned limitations, we first dive deeper into the discriminative feature fusion strategies presented in recent universal semantic segmentation studies [1, 7]. Subsequently, we introduce a novel heterogeneous feature fusion block (HFB) to process the RGB and surface normal features, which are encoded using two independent Swin Transformer backbones [10]. Our technical contributions in this part are three-fold: (1) a holistic attention module (HAM) to model the interdependencies between heterogeneous features across three dimensions (spatial, channel, and scale); (2) a heterogeneous feature contrast descriptor (HFCD) to effectively underscore both the shared and unique characteristics within the holistically-attentive features; (3) an affinity-weighted feature recalibrator (AWFR) to jointly emphasize and suppress heterogeneous features prior to their input into the decoder. Additionally, we contribute to a lightweight yet more effective decoder, which incorporates inter-scale skip connections while pruning redundant ones. Our decoder demonstrates greater feature decoding capabilities, while simultaneously reducing computational complexity. Finally, we design two new loss functions based on semantic annotations and depth data to provide deeper supervision during our model training process. This contribution also results in improved overall performance, particularly in error-prone areas. The effectiveness of each contribution is validated through extensive experiments conducted across public datasets. Excitingly, our proposed freespace detection framework SNE-RoadSegV2, as shown in Fig. 1, which incorporates all these innovative components, wins the 1st place on the KITTI Road leaderboard [11]. In a nutshell, our contributions are as follows:
1.
We propose SNE-RoadSegV2, a novel feature-fusion freespace detection approach, achieving state-of-the-art (SoTA) performance across multiple public datasets. Our source code will be released upon publication.
2.
We introduce HFB, consisting of an HAM, an HFCD, and an AWFR, for comprehensive heterogeneous feature description, recalibration, and fusion, resulting in more coherent feature representations.
3.
We design a lightweight yet more effective decoder, incorporating inter-scale skip connections while pruning redundant intra-scale ones, demonstrating greater efficacy and computational efficiency.
4.
We develop two novel fallibility-aware loss functions, which focus particularly on reducing misclassifications in semantic-transition and depth-inconsistent regions, leading to improved overall performance.
II Related Work
II-A Data-Driven Freespace Detection
While it is feasible to employ universal semantic segmentation networks [12, 13, 14, 15] for this task, it has been observed that task-specific approaches [16, 1, 7] consistently deliver superior performance. Early task-specific freespace detection approaches [17, 18, 19, 20] thoroughly rely on RGB images and were found to be highly sensitive to environmental factors, notably illumination and weather conditions [1]. Given the increased prevalence of range sensors, particularly LiDARs, feature-fusion networks [2, 21, 22] have emerged as a more robust choice in this domain. In terms of network architecture, these approaches are characterized by duplex-encoder architectures [1, 16], where each encoder extracts hierarchical features from a specific data source or modality. The extracted heterogeneous features are subsequently fused, enabling the network to gain a more comprehensive understanding of the environment [2]. As for the input data, the most commonly used spatial geometric information includes depth/disparity maps [23, 24], LiDAR point clouds [2, 21], and surface normal information [1, 7, 16]. Extensive experiments conducted in previous studies [1, 7, 16] have conclusively demonstrated that surface normals provide the most informative spatial geometric information for freespace detection, owing to their representation of planar characteristics. Therefore, in this paper, we adopt the pipeline introduced in [1, 7, 16], which utilizes a duplex-encoder architecture to extract heterogeneous features from RGB images and surface normal information. However, it is important to note that our focus differs from these previous work. Our emphasis lies in designing heterogeneous feature fusion strategies, developing a lightweight yet more effective decoder, and introducing task-specific loss functions.
II-B Heterogeneous Feature Fusion
Heterogeneous feature fusion plays a pivotal role in various computer vision tasks, such as salient object detection [25, 26, 27] and scene parsing [28, 29, 30]. Although this topic may not have received extensive attention in freespace detection, it is worth noting that there have been notable prior works proposed in the broader field of semantic segmentation. For instance, Wei et al. [25] proposed the cross feature module (CFM), capable of refining features at multiple levels while simultaneously suppressing background noise. Additionally, the separation-and-aggregation gate (SAGate) [26] is another general-purpose heterogeneous feature fusion method that incorporates complementary information through feature recalibration and aggregation to generate selective representations for segmentation. Moreover, Pang et al. [26] introduced the dynamic dilated pyramid module (DDPM), which generates adaptive kernels for efficient feature decoding. As these prior studies generally overlook the appropriate discrimination between the inherent differences between heterogeneous features, our primary focus in this paper is directed towards addressing this aspect.
II-C Attention Mechanisms
Attention mechanisms are vital components within modern deep learning models, allowing for effective concentration on specific elements of input data, ultimately leading to a more comprehensive understanding of the environment [31, 32, 33]. As a representative example, the squeeze-and-excitation network (SENet) [31] dynamically recalibrates channel-wise feature responses by explicitly modeling dependencies between channels, enabling the network to emphasize informative channels while suppressing less relevant ones. In addition to channel attention, the convolutional block attention module (CBAM) [34] introduces attention from another dimension – spatial. This lightweight and highly compatible module sequentially computes channel and spatial attention maps and multiplies them with the input feature maps to achieve more adaptive feature refinement. On the other hand, Swin Transformer [10] is a general-purpose Transformer backbone developed specifically for fundamental computer vision tasks. It is highly regarded for its hierarchical representation learning approach, which computes self-attention locally within non-overlapping shifted windows. This innovative design contributes to its compelling performance in applications including image classification, object detection, and semantic segmentation. In this paper, we first extend CBAM to three dimensions: the original two plus scale. Moreover, we use two Swin Transformers as the backbone networks in our duplex encoder, and comprehensive experiments in the supplement provide evidence of its superior performance compared to CNNs.
III Methodology
III-A Architecture Overview
Fig. 1 provides readers with an overview of the SNE-RoadSegV2 architecture, consisting of three key elements: duplex feature embedding, heterogeneous feature fusion, and lightweight yet effective feature decoding. A pair of input RGB image and surface normal map , translated from a depth image using a surface normal estimator (SNE) [1], are first tokenized into non-overlapping patches and transformed into a high-dimensional feature space through a trainable linear projection in the patch embedding module [10]. The embedded features are subsequently fed into the Swin Transformer blocks [10] to produce hierarchical heterogeneous features and . Each pair of heterogeneous features undergoes a comprehensive fusion process through HFB. Finally, a lightweight yet more effective decoder which incorporates both inter-scale and intra-scale skip connections is designed to further boost the efficiency and accuracy of freespace detection. The proposed architecture is trained by minimizing a loss with fallibility-awareness incorporated. The following subsections will provide a detailed explanation of the HFB, decoder, and loss functions in sequence.
III-B Heterogeneous Feature Fusion Block
The core problem of feature encoding for freespace detection revolves around the effective fusion of heterogeneous features extracted from various data sources. As detailed in Sec. I, prior investigations, e.g. , [1, 7, 23, 21] fused these heterogeneous features without appropriate discrimination between their inherent differences. Our HFB is, thus, designed to overcome this limitation. As depicted in Fig. 2, meaningful heterogeneous features are first selectively emphasized and suppressed across three dimensions (spatial, channel, and scale) through an HAM. These features are further enhanced through an HFCD, which improves the representations of their shared and distinct characteristics. The original heterogeneous features are ultimately weighted through an AWFR to emphasize the aspects that are important to both or either of the features.
III-B1 Holistic Attention Module
Before contrasting and fusing heterogeneous features, it is imperative to emphasize or attenuate specific spatial regions and channels across multiple scales [34, 15]. Drawing inspiration from the CBAM [34], we first apply spatial attention to both RGB and surface normal feature maps and , resulting in spatially weighted feature maps and , respectively, which are then concatenated to form :
where denotes the spatial attention operation, allowing the model to prioritize important regions of an image while suppressing less relevant areas.
As illustrated in Fig. 2, another branch of HAM incorporates multi-scale contextual information along with channel attention. A series of atrous convolutional layers [15] are initially employed to generate features with progressively expanded receptive fields:
where denotes the multi-scale context aggregation operation. A channel attention operation is subsequently applied to further model the interdependencies between heterogeneous features at both scale and channel levels, resulting in scale-channel attentive feature maps as follows:
The expressions of , , and are detailed in the supplement. In contrast to prior arts [34, 31] which exclusively focus on channel attention and overlook the consideration of multiple scales, our designed scale-channel co-attention mechanism provides a more comprehensive perspective on heterogeneous features.
III-B2 Heterogeneous Feature Contrast Descriptor
After modeling the interdependencies of the heterogeneous features across three separate dimensions using (1)-(3), we further explore the way to contrast these features in a more comprehensive and effective fashion. Unlike relevant prior arts, e.g. , CFM [25], SAGate [26], and DDPM [27], which primarily emphasize feature commonality, our investigation delves deeper into both their shared and distinct characteristics. To this end, we first normalize the spatial-attentive and scale-channel co-attentive features and using a sigmoid function , yielding and , respectively. Performing an element-wise product operation between and activates jointly-emphasized parts between heterogeneous features, while performing an element-wise subtraction operation between and activates features that are important only in either the RGB image or the surface normal map. Upon such inspiration, we formulate our HFCD as follows:
where refers to the Hadamard product, denotes element-wise subtraction, and (with its expression provided in the supplement) represents a combination of convolutional, batchnorm, and sigmoid layers. As shown in Fig. 2, the described heterogeneous feature contrast is subsequently utilized to construct an affinity volume for further feature recalibration.
III-B3 Affinity-Weighted Feature Recalibrator
Constructing a volume that contains element-wise weights to jointly recalibrate (emphasize and de-emphasize) heterogeneous features is another significant contribution in our designed HFB. As and describe the contrasting aspects between the heterogeneous features at the spatial and scale-channel levels, respectively, we employ these two volumes to construct an affinity volume as follows:
which provides the original heterogeneous features and with element-wise weights between 0 and 1. Specifically, a higher affinity value indicates a greater importance of that element in both types of feature maps, an intermediate affinity value indicates a greater importance of element in either of the feature maps, while a lower affinity value indicates a redundant element in both types of feature maps that should be neglected. Finally, and are weighted by to form and for the next stage of feature encoding, and then concatenated to generate the recalibrated heterogeneous features as follows:
where , , and denote convolutional layers. Compared to other SoTA heterogeneous feature fusion strategies, our proposed HFB adaptively assigns weights to the original features, taking into account both the elements of agreement and disagreement in importance as determined by HFCD and AWFR. Such a way of seeking common ground and while preserving differences enhances the comprehensiveness of feature fusion for freespace detection. The superior performance of HFB and the effectiveness of its each component are demonstrated in Sec. IV-D.
III-C Lightweight yet More Effective Decoder
Fig. 3 illustrates the decoder architectures of RoadSeg [1] (identical to UNet++ [35]), UNet3+ [9], and our proposed SNE-RoadSegV2. As claimed in [9], UNet++ fails to effectively utilize multi-scale features, and UNet3+ was designed specifically to resolve this limitation. However, before designing our decoder, we must address the question: does UNet3+ consistently outperform UNet++ in the freespace detection task? Through an extensive series of experiments with both CNN and Transformer backbones, as detailed in the supplement, we regret to report that the answer is negative. It is our contention that both the intra-scale skip connections in UNet++ and the inter-scale skip connections in UNet3+ remain indispensable. Thus, we design the SNE-RoadSegV2 decoder, which combines the strengths of both UNet++ and UNet3+, through enormous experimental efforts. As shown in Fig. 3, we maintain skip connections only from a given node to its adjacent and final nodes at each stage. This modification reduces redundant information propagation without compromising decoder performance. Additionally, we adopt the inter-scale skip connections used in UNet3+ to capture both fine-grained and coarse-grained details. Moreover, we replace the basic convolutions in the decoder with depth-wise separable convolutions [36] to further reduce its computational complexity. Sec. IV-D quantitatively demonstrates that our SNE-RoadSegV2 decoder outperforms UNet++ and UNet3+ in terms of both efficiency and accuracy. Despite its superior performance, we consider the decoder as an experimental contribution.
III-D Fallibility-Aware Loss Functions
The pixel-wise BCE loss has been a common choice in previous studies as the primary criterion to supervise model training. Nonetheless, it is unfortunate that these previous efforts have not considered the specific characteristics of real-world driving scenarios. Misclassifications are frequently observed near the transition regions between different semantic categories [25]. In addition, the depth data are not explicitly utilized to supervise model training. Hence, we propose a novel modification to the conventional BCE loss function by introducing two weighting factors that prioritize these error-prone regions. The overall loss is formulated as follows:
where and balance the semantic transition-aware loss and the depth inconsistency-aware loss , respectively. The ablation studies on their individual efficacy and the selection of hyper-parameters and are provided in Sec. IV-D.
III-D1 Semantic Transition-Aware Loss
A transition region between different semantic categories can be considered as a mixture of semantic labels. For each pixel with a neighborhood system , where and denote the foreground (freespace) and background (others) sets, respectively, we determine its likelihood of belonging to a semantic transition region using the following expression:
where is the indicator function, and . (8) approaches 0 when either or is close to being an empty set, and approaches 1 vice versa. Substituting (8) into the standard BCE loss yields as follows:
where denotes the ground-truth label of (1 for freespace, and 0 otherwise), while indicates the probability that belongs to the freespace category. Our supplement provides details on the selection of radius.
III-D2 Depth Inconsistency-Aware Loss
When depth images are available, it is also advantageous to leverage these data to improve network training via an adaptive loss function. Surprisingly, prior research efforts have not explored this aspect.
Let be a matrix storing the homogeneous coordinates of the predicted freespace pixels . Considering that the height between the camera and the ground plane remains theoretically constant in each image, when disregarding the camera pitch angle, we aggregate the -coordinates of these pixels to derive a theoretical camera height, denoted as , using the following expression:
where stores the depth values . A weight measuring the depth inconsistency can then be yielded as follows:
which approaches 1 when depth is inconsistent, and approaches 0 vice versa. is thus formulated as follows:
| Method | Input Data | MaxF (%) | AP (%) | Pre (%) | Rec (%) | FPR (%) | FNR (%) | Rank |
| Hadamard-FCN [38] | RGB | 94.85 | 91.48 | 94.81 | 94.89 | 2.86 | 5.11 | 29 |
| RBANet [24] | RGB | 96.30 | 89.72 | 95.14 | 97.50 | 2.75 | 2.50 | 22 |
| HA-DeepLabv3+ [39] | RGB + Disp | 94.83 | 93.24 | 94.77 | 94.89 | 2.88 | 5.11 | 30 |
| DFM-RTFNet [40] | RGB + Disp | 96.78 | 94.05 | 96.62 | 96.93 | 1.87 | 3.07 | 11 |
| USNet [23] | RGB + Depth | 96.89 | 93.25 | 96.51 | 97.27 | 1.94 | 2.73 | 9 |
| LRDNet+ [21] | RGB + LiDAR PC | 96.95 | 92.22 | 96.88 | 97.02 | 1.72 | 2.98 | 8 |
| PLB-RD [41] | RGB + LiDAR PC | 97.42 | 94.09 | 97.30 | 97.54 | 1.49 | 2.46 | 4 |
| PLARD [2] | RGB + LiDAR PC | 97.03 | 94.03 | 97.19 | 96.88 | 1.54 | 3.12 | 7 |
| BJN [42] | RGB + LiDAR PC | 94.89 | 90.63 | 96.14 | 93.67 | 2.07 | 6.33 | 26 |
| LidCamNet [30] | RGB + LiDAR PC | 96.03 | 93.93 | 96.23 | 95.83 | 2.07 | 4.17 | 18 |
| CLCFNet [22] | RGB + LiDAR PC | 96.38 | 90.85 | 96.38 | 96.39 | 1.99 | 3.61 | 15 |
| NIM-RTFNet [43] | RGB + Normal | 96.02 | 94.01 | 96.43 | 95.62 | 1.95 | 4.38 | 19 |
| SNE-RoadSeg [1] | RGB + Normal | 96.75 | 94.07 | 96.90 | 96.61 | 1.70 | 3.39 | 12 |
| SNE-RoadSeg+ [7] | RGB + Normal | 97.50 | 93.98 | 97.41 | 97.58 | 1.43 | 2.42 | 3 |
| RoadFormer [16] | RGB + Normal | 97.50 | 93.85 | 97.16 | 97.84 | 1.57 | 2.16 | 2 |
| SNE-RoadSegV2 (Ours) | RGB + Normal | 97.55 | 93.98 | 97.57 | 97.53 | 1.34 | 2.47 | 1 |
| Method | Fsc (%) | IoU (%) | Acc (%) |
| NIM-RTFNet [43] | 92.02 | 91.43 | 96.07 |
| RBANet [24] | 93.81 | 91.88 | 96.50 |
| USNet [23] | 94.28 | 92.74 | 96.71 |
| LRDNet+ [21] | 94.71 | 92.82 | 97.02 |
| PLARD [2] | 95.28 | 92.96 | 97.15 |
| SNE-RoadSeg [1] | 96.49 | 93.22 | 97.68 |
| SNE-RoadSegV2 (Ours) | 97.12 | 94.40 | 98.11 |
IV Experiments
First, we conduct extensive experiments on the KITTI road dataset [11] (medium-sized) and the Cityscapes dataset [44] (large-scale) to validate the effectiveness of our proposed encoder, decoder, and loss functions. Subsequently, we perform both qualitative and quantitative comparisons between SNE-RoadSegV2 and other SoTA freespace detection algorithms. To further demonstrate the superior performance of our network, we also provide additional experiments carried out on the vKITTI2 dataset [45] (large-scale yet synthetic) and the KITTI semantics dataset [46] (real-world yet small-sized) to validate the robustness of our freespace detection framework.
IV-A Datasets and Evaluation Metrics
The details on the aforementioned four datasets are as follows:
KITTI Road [11]: this dataset provides real-world RGB-D data (image resolution: 1,242375 pixels) for the evaluation of data-fusion freespace detection algorithms. Following the study presented in [1], we split the dataset into three subsets: training (173 images), validation (58 images), and test (58 images) to conduct ablation studies and hyper-parameter selection experiments.
Cityscapes [44]: this dataset provides real-world stereo images (resolution: 2,0481,024 pixels), each manually annotated with 34 semantic classes. In our experiments, we preprocess the ground-truth annotations by categorizing them into two groups: freespace and others. As there is no depth ground truth is available, we utilize a pre-trained RAFT-Stereo [47] to generate depth images. We adhere to the official split of training, and validation sets, with 2,975 and 500 images in each set.
KITTI Semantics [11]: this dataset contains 200 real-world RGB images, each accompanied by well-annotated semantic ground truth for 19 different classes (in alignment with the Cityscapes [44] dataset). In our experiments, we pre-process the ground-truth annotations by categorizing them into two groups: freespace and others. Sparse disparity ground truth is obtained using a Velodyne HDL-64E LiDAR. We generate dense depth maps using a well-trained CreStereo [48] model in our experiments. These images are randomly divided into a training set and a validation set, with a ratio of 3:1.
vKITTI2 [45]: this dataset contains virtual replicas of five sequences from the KITTI dataset. Dense ground-truth depth maps are acquired through depth rendering using a virtual engine. In our experiments, we use the first four sequences for model training, and validate the model performance on the remaining sequence.
Adhering to the experiments presented in [1], we quantify the model’s performance using accuracy (Acc), precision (Pre), recall (Rec), F1-score (Fsc), and intersection over union (IoU). Additionally, when submitting the results obtained from the best-performing model to the KITTI Road benchmark, we also compute maximum F1-measure (MaxF), average precision (AP), false-positive rate (FPR), and false negative rate (FNR).
IV-B Implementation Details
Our experiments are conducted using an Intel Core i7-12700k CPU and an NVIDIA RTX 4090 GPU. The Adam optimizer [49] with an initial learning rate of 0.001 is used to minimize the loss function. A multi-step learning scheduler with a decay rate of 0.5 for every 20 epochs is also employed. Each model is trained for a total of 100 epochs, with early stopping mechanisms applied to the validation set to prevent over-fitting. Common data augmentation techniques, such random flipping, rotation, cropping, and brightness adjustment, are also applied to enhance the model’s robustness.
| Method | Fsc (%) | IoU (%) | Acc (%) |
| NIM-RTFNet [43] | 92.59 | 85.95 | 96.61 |
| OFF-Net [50] | 93.82 | 86.79 | 97.08 |
| SNE-RoadSeg [1] | 94.85 | 88.02 | 97.33 |
| SNE-RoadSeg+ [7] | 95.11 | 89.07 | 97.59 |
| RoadFormer [16] | 95.36 | 90.18 | 97.83 |
| SNE-RoadSegV2 (Ours) | 96.60 | 91.75 | 98.44 |
| Method | Fsc (%) | IoU (%) | Acc (%) |
| NIM-RTFNet [43] | 92.39 | 85.85 | 96.91 |
| OFF-Net [50] | 93.61 | 87.53 | 97.22 |
| SNE-RoadSeg [1] | 94.03 | 88.74 | 97.63 |
| SNE-RoadSeg+ [7] | 95.20 | 90.50 | 97.93 |
| RoadFormer [16] | 95.39 | 91.72 | 98.31 |
| SNE-RoadSegV2 (Ours) | 96.47 | 92.33 | 98.93 |
IV-C Comparison with State-of-the-art Methods
The quantitative and qualitative experimental results on the KITTI Road dataset are presented in Table I and Fig. 4, respectively, while the quantitative and qualitative experimental results on the Cityscapes dataset are given in Table II and Fig. 5, respectively. These results suggest that our proposed SNE-RoadSegV2 demonstrates superior performance compared to all other SoTA freespace detection approaches, with an increase in MaxF up to 2.72% on the KITTI dataset and an increase in IoU by 1.18-2.97% on the Cityscapes dataset. Notably, it achieves the 1st rank on the KITTI Road benchmark. The qualitative comparisons, with significantly improved regions highlighted by red dashed boxes, particularly near semantic-transition and depth-inconsistent regions, also validate the effectiveness of our designed feature fusion block, decoder, and loss function. The ablation studies that validate the individual efficacy of these components are discussed in the next subsection.
The quantitative comparisons among SoTA freespace detection algorithms on the KITTI Semantics [11] and VKITTI2 datasets are presented in Tables III and IV, respectively, while their qualitative comparisons on these two datasets are provided in Figs. 7 and 8. These results suggest that our proposed SNE-RoadSegV2 outperforms all other SoTA approaches, with an increase in Fsc by 1.24-4.01% on the KITTI Semantics dataset, and by 1.08-4.08% on the VKITTI2 dataset. It is noteworthy that our SNE-RoadSegV2 demonstrates greater robustness, particularly near/on semantic transition regions or in areas with surface normals similar to that of the road surface. The achieved improvements are primarily attributed to our heterogeneous feature fusion strategy and fallibility-aware loss functions.
We also present experimental results across various road scene categories: (1) urban marked (UM), (2) urban multiple marked (UMM), and (3) urban unmarked (UU), as specified in the KITTI Road benchmark at cvlibs.net/datasets/kitti/eval_road.php. As shown in Table V, our proposed SNE-RoadSegV2 achieves SoTA performance in both UM and UU scenes, and demonstrates comparable performance in the UMM scene, when compared to the recent SoTA algorithm, RoadFormer [16]. We also provide illustrative experimental results in the bird’s eye view (BEV) in Fig. 9. These results clearly demonstrate the robust performance of our SNE-RoadSegV2 in detecting freespace.
| Methods | Road Scenes | ||
| UM | UMM | UU | |
| CLCFNet [22] | 95.65 | 97.24 | 95.68 |
| NIM-RTFNet [43] | 95.71 | 96.79 | 95.11 |
| RBANet [24] | 95.78 | 97.38 | 94.91 |
| LRDNet+ [21] | 96.10 | 97.98 | 96.18 |
| SNE-RoadSeg [1] | 96.42 | 97.47 | 96.03 |
| DFM-RTFNet [40] | 96.46 | 97.45 | 96.26 |
| USNet [23] | 96.46 | 97.68 | 96.11 |
| PLB-RD [41] | 96.87 | 98.05 | 95.63 |
| SNE-RoadSeg+ [7] | 96.95 | 98.13 | 97.04 |
| RoadFormer [16] | 97.02 | 98.15 | 97.02 |
| PLARD [2] | 97.05 | 97.77 | 95.95 |
| SNE-RoadSegV2 (Ours) | 97.25 | 98.10 | 97.08 |
| Backbone | Params (M) | KITTI Road Dataset | Cityscapes Dataset | ||
| Fsc (%) | IoU (%) | Fsc (%) | IoU (%) | ||
| Eff-B0 | 11.85 | 95.98 | 92.36 | 96.59 | 95.57 |
| Eff-B1 | 17.22 | 96.08 | 92.46 | 97.14 | 95.05 |
| Eff-B2 | 20.29 | 96.33 | 92.92 | 95.68 | 96.99 |
| Eff-B3 | 28.00 | 96.44 | 93.12 | 95.85 | 97.03 |
| Eff-B4 | 44.78 | 96.78 | 93.75 | 97.10 | 96.46 |
| Eff-B5 | 69.46 | 96.21 | 92.69 | 97.38 | 95.06 |
| Eff-B6 | 98.83 | 96.26 | 92.79 | 96.47 | 96.05 |
| Eff-B7 | 151.80 | 96.44 | 93.12 | 96.38 | 96.49 |
| Res-18 | 34.31 | 95.87 | 92.07 | 96.88 | 94.89 |
| Res-34 | 59.92 | 95.86 | 92.04 | 96.84 | 94.89 |
| Res-50 | 188.75 | 96.27 | 92.81 | 96.92 | 95.63 |
| Res-101 | 245.73 | 96.47 | 93.18 | 96.76 | 96.19 |
| Res-152 | 292.66 | 96.92 | 94.02 | 96.92 | 96.92 |
| Swin-T | 66.53 | 96.34 | 92.94 | 96.62 | 96.07 |
| Swin-S | 109.12 | 96.91 | 93.26 | 97.11 | 97.02 |
| Swin-B | 193.85 | 97.44 | 94.13 | 97.25 | 97.24 |
IV-D Ablation Study
| Decoder | Fsc (%) | IoU (%) | Params (M) | FLOPS (G) |
| UNet++ [35] / RoadSeg [1] | 96.27 | 93.90 | 13.62 | 78.44 |
| UNet3+ [9] | 95.64 | 93.41 | 14.70 | 164.63 |
| SNE-RoadSegV2 (Ours) [9] | 97.58 | 94.50 | 6.71 | 60.33 |
IV-D1 Selection of Duplex Encoder Backbone
We evaluate the heterogeneous feature extraction performance using various backbones, including EfficientNet [51], ResNet [52], and Swin Transformer [10]. As shown in Table VI, the backbones based on Swin Transformer outperform those based on EfficientNet and ResNet in terms of Fsc and IoU. Notably, the Swin-B backbone achieves the best overall performance among all the backbones on both the KITTI Road and Cityscapes datasets. Consequently, we select Swin-B as the network backbone for our model.
IV-D2 Encoder
We first explore the rationality of each component in HFB. As presented in Table VIII, we sequentially remove each component from HFB to quantify its impact on the overall performance. It is evident that each component in HFB contributes to an improvement in the overall performance, and the network achieves its peak performance when all components (HAM, HFCD, and AWFR) are integrated into HFB, which demonstrates the effectiveness of our design.
Additionally, we compare HFB with other SoTA heterogeneous feature fusion strategies. As shown in Table IX, HFB outperforms other compared methods on both datasets, with improvements of up to 1.26% in Fsc and 1.67% in IoU, respectively. These compelling results can be attributed to our novel contributions, particularly the exploitation of both shared and distinct characteristics of heterogeneous features in HFCD, and the effective feature recalibration based on the affinity volume constructed in AWFR. Another qualitative comparison illustrated in Fig. 6 further validates the effectiveness of HFB. Obviously, HFB produces more coherent feature representations with clearer segmentation boundaries, compared to the indiscriminate feature fusion strategy adopted in SNE-RoadSeg.
| Baseline | HAM | HFCD | ACFR | Fsc (%) | ||
| SA | CA | AC | ||||
| ✓ | 96.65 | |||||
| ✓ | ✓ | ✓ | 96.54 | |||
| ✓ | ✓ | ✓ | ✓ | 97.11 | ||
| ✓ | ✓ | ✓ | ✓ | ✓ | 97.69 | |
| ✓ | ✓ | ✓ | ✓ | 96.84 | ||
| ✓ | ✓ | ✓ | ✓ | 97.08 | ||
| Strategy | KITTI Road Dataset | Cityscapes Dataset | ||
| Fsc (%) | IoU (%) | Fsc (%) | IoU (%) | |
| CFM [25] | 95.15 | 91.58 | 95.45 | 92.33 |
| SAGate [26] | 96.77 | 93.74 | 96.80 | 93.81 |
| DDPM [27] | 96.05 | 92.98 | 96.60 | 93.43 |
| HFB (Ours) | 97.68 | 94.90 | 97.12 | 94.40 |
IV-D3 Decoder
Quantitative comparisons of decoder performance among SNE-RoadSegV2, RoadSeg/UNet++, and UNet3+ are presented in Table. VII. These results demonstrate the SoTA performance of our decoder, with improvements in Fsc and IoU by up to 1.94% and 1.09%, respectively, while maintaining lower computational complexity, including a reduction in learnable parameters and FLOPS by up to 54.35% and 63.35%, respectively. These improvements can be attributed to the use of depth-wise separable convolution and the pruning of redundant skip connections.
IV-D4 Loss Function
Fig. 10 actually presents two experiments: (1) when (only conventional BCE loss is used), the overall freespace detection performance on both datasets is the worst, demonstrating the effectiveness of our proposed fallibility-aware losses; (2) different ratios between and demonstrate that when and (the sum of these two weights is empirically set to 0.4, as discussed in [53, 54, 55]), SNE-RoadSegV2 achieves the best performance on both datasets. While further hyper-parameter tuning is possible, it is important to consider the risk of over-fitting with limited data.
As discussed in Sec. 3.4.1, in order to maximize the effectiveness of the proposed semantic transition-aware loss , we conduct experiments with a series of neighborhood system radii. As illustrated in Fig. 11, SNE-RoadSegV2 achieves the best overall performance on the KITTI Road dataset when the neighborhood radius is set to pixels.
V Conclusion
This paper revisited the designs of heterogeneous feature fusion strategies, decoder architectures, and loss functions from prior research, and introduced SNE-RoadSegV2, a novel, high-performing, state-of-the-art freespace detection network. Breaking down our contributions further, our work contains six technical contributions: three novel components in the encoder, one decoder architecture, and two loss functions. The effectiveness of each contribution was validated through extensive experiments. Comprehensive comparisons with other state-of-the-art algorithms unequivocally demonstrate the superiority of SNE-RoadSegV2, notably achieving the top rank on the KITTI Road benchmark. Our future work will mainly focus on extending SNE-RoadSegV2 to other general-purpose semantic segmentation tasks.

### v2

@@
 To address the aforementioned limitations, we first dive deeper into the discriminative feature fusion strategies presented in recent universal semantic segmentation studies [1, 7].
-Subsequently, we introduce a novel heterogeneous feature fusion block (HFB) to process the RGB and surface normal features, which are encoded using two independent Swin Transformer backbones [10].
+Subsequently, we introduce a novel heterogeneous feature fusion block (HF2B) to process the RGB and surface normal features, which are encoded using two independent Swin Transformer backbones [10].
 Our technical contributions in this part are three-fold: (1) a holistic attention module (HAM) to model the interdependencies between heterogeneous features across three dimensions (spatial, channel, and scale); (2) a heterogeneous feature contrast descriptor (HFCD) to effectively underscore both the shared and unique characteristics within the holistically-attentive features; (3) an affinity-weighted feature recalibrator (AWFR) to jointly emphasize and suppress heterogeneous features prior to their input into the decoder.
@@
 2.
-We introduce HFB, consisting of an HAM, an HFCD, and an AWFR, for comprehensive heterogeneous feature description, recalibration, and fusion, resulting in more coherent feature representations.
+We introduce HF2B, consisting of an HAM, an HFCD, and an AWFR, for comprehensive heterogeneous feature description, recalibration, and fusion, resulting in more coherent feature representations.
 3.
@@
 The embedded features are subsequently fed into the Swin Transformer blocks [10] to produce hierarchical heterogeneous features and .
-Each pair of heterogeneous features undergoes a comprehensive fusion process through HFB.
+Each pair of heterogeneous features undergoes a comprehensive fusion process through HF2B.
 Finally, a lightweight yet more effective decoder which incorporates both inter-scale and intra-scale skip connections is designed to further boost the efficiency and accuracy of freespace detection.
 The proposed architecture is trained by minimizing a loss with fallibility-awareness incorporated.
-The following subsections will provide a detailed explanation of the HFB, decoder, and loss functions in sequence.
+The following subsections will provide a detailed explanation of the HF2B, decoder, and loss functions in sequence.
 III-B Heterogeneous Feature Fusion Block
@@
 , [1, 7, 23, 21] fused these heterogeneous features without appropriate discrimination between their inherent differences.
-Our HFB is, thus, designed to overcome this limitation.
+Our HF2B is, thus, designed to overcome this limitation.
 As depicted in Fig.
@@
 III-B3 Affinity-Weighted Feature Recalibrator
-Constructing a volume that contains element-wise weights to jointly recalibrate (emphasize and de-emphasize) heterogeneous features is another significant contribution in our designed HFB.
+Constructing a volume that contains element-wise weights to jointly recalibrate (emphasize and de-emphasize) heterogeneous features is another significant contribution in our designed HF2B.
 As and describe the contrasting aspects between the heterogeneous features at the spatial and scale-channel levels, respectively, we employ these two volumes to construct an affinity volume as follows:
@@
 where , , and denote convolutional layers.
-Compared to other SoTA heterogeneous feature fusion strategies, our proposed HFB adaptively assigns weights to the original features, taking into account both the elements of agreement and disagreement in importance as determined by HFCD and AWFR.
+Compared to other SoTA heterogeneous feature fusion strategies, our proposed HF2B adaptively assigns weights to the original features, taking into account both the elements of agreement and disagreement in importance as determined by HFCD and AWFR.
 Such a way of seeking common ground and while preserving differences enhances the comprehensiveness of feature fusion for freespace detection.
-The superior performance of HFB and the effectiveness of its each component are demonstrated in Sec.
+The superior performance of HF2B and the effectiveness of its each component are demonstrated in Sec.
 IV-D.
@@
 The quantitative comparisons among SoTA freespace detection algorithms on the KITTI Semantics [11] and VKITTI2 datasets are presented in Tables III and IV, respectively, while their qualitative comparisons on these two datasets are provided in Figs.
-7 and 8.
+9 and 10.
 These results suggest that our proposed SNE-RoadSegV2 outperforms all other SoTA approaches, with an increase in Fsc by 1.24-4.01% on the KITTI Semantics dataset, and by 1.08-4.08% on the VKITTI2 dataset.
@@
 We also provide illustrative experimental results in the bird’s eye view (BEV) in Fig.
-9.
+11.
 These results clearly demonstrate the robust performance of our SNE-RoadSegV2 in detecting freespace.
@@
 IV-D2 Encoder
-We first explore the rationality of each component in HFB.
-As presented in Table VIII, we sequentially remove each component from HFB to quantify its impact on the overall performance.
-It is evident that each component in HFB contributes to an improvement in the overall performance, and the network achieves its peak performance when all components (HAM, HFCD, and AWFR) are integrated into HFB, which demonstrates the effectiveness of our design.
-Additionally, we compare HFB with other SoTA heterogeneous feature fusion strategies.
-As shown in Table IX, HFB outperforms other compared methods on both datasets, with improvements of up to 1.26% in Fsc and 1.67% in IoU, respectively.
+We first explore the rationality of each component in HF2B.
+As presented in Table VIII, we sequentially remove each component from HF2B to quantify its impact on the overall performance.
+It is evident that each component in HF2B contributes to an improvement in the overall performance, and the network achieves its peak performance when all components (HAM, HFCD, and AWFR) are integrated into HF2B, which demonstrates the effectiveness of our design.
+Additionally, we compare HF2B with other SoTA heterogeneous feature fusion strategies.
+As shown in Table IX, HF2B outperforms other compared methods on both datasets, with improvements of up to 1.26% in Fsc and 1.67% in IoU, respectively.
 These compelling results can be attributed to our novel contributions, particularly the exploitation of both shared and distinct characteristics of heterogeneous features in HFCD, and the effective feature recalibration based on the affinity volume constructed in AWFR.
 Another qualitative comparison illustrated in Fig.
-6 further validates the effectiveness of HFB.
-Obviously, HFB produces more coherent feature representations with clearer segmentation boundaries, compared to the indiscriminate feature fusion strategy adopted in SNE-RoadSeg.
+6 further validates the effectiveness of HF2B.
+Obviously, HF2B produces more coherent feature representations with clearer segmentation boundaries, compared to the indiscriminate feature fusion strategy adopted in SNE-RoadSeg.
 | Baseline | HAM | HFCD | ACFR | Fsc (%) | ||
@@
 | DDPM [27] | 96.05 | 92.98 | 96.60 | 93.43 |
-| HFB (Ours) | 97.68 | 94.90 | 97.12 | 94.40 |
+| HF2B (Ours) | 97.68 | 94.90 | 97.12 | 94.40 |
 IV-D3 Decoder
@@
 As illustrated in Fig.
-11, SNE-RoadSegV2 achieves the best overall performance on the KITTI Road dataset when the neighborhood radius is set to pixels.
+13, SNE-RoadSegV2 achieves the best overall performance on the KITTI Road dataset when the neighborhood radius is set to pixels.
 V Conclusion